Calculate the pH of black coffee with a hydrogen ion concentration of 4.55 × 10−6 mol dm−3. Show all steps of your calculations and include the correct units where needed. Give your pH value to three significant figures. Is coffee an acidic or basic solution? Explain your choice.


1
Expert's answer
2022-02-10T20:07:02-0500

Using C = 4.55 × 10−6 mol dm−3 for the H+ concentration, we determine the pH as


"pH=-\log{[H^+]}=- \log{(\frac{4.55 × 10^{−6}\,\cancel{M}}{1\,\cancel{M}})}
\\ pH= -\log{(4.55 × 10^{−6})}
\\ pH=-\log{(4.55)}-\log{(10^{−6})}
\\ pH= -0.66-({−6})=6-0.66
\\ \therefore pH = 5.34"


In conclusion, the pH of black coffee is 5.34, since pH<7 coffee is an acidic solution.